# Joint Estimation of Identity Verification and Relative Pose for Partial Fingerprints

### v1

@@ Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. Consequently, in this paper we propose a method that jointly estimates identity verification and relative pose for partial fingerprints, aiming to leverage their inherent correlation to improve each other. To achieve this, we propose a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset show that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods.
@@ I Introduction
Fingerprints are unique patterns composed of ridges and valleys on the finger surface. Due to its easy collection, high stability and strong recognizability, this biological characteristic exhibits significant application value. As early as the nineteenth century, researchers had conducted systematic studies on fingerprints [1, 2, 3]. The reliability of fingerprint recognition gained official recognition in the twentieth century and found extensive application within the judicial sphere [3]. With the advancement of devices and algorithms, fingerprint recognition technology has been swiftly applied in civilian and commercial fields, such as electronic devices, financial services, and access control systems, providing people with convenience and security [4].
@@
Fueled by consumer demand and advances in technology, the trend towards integration, miniaturization, and portability has become prominent within the consumer electronics sector. This progression compels manufacturers to consistently reduce the dimensions of sensors to accommodate progressively compact device architectures. Fingerprint sensors, currently one of the most popular biometric modules, have also introduced various miniaturized solutions for extensive use [4]. However, the size of fingerprint sensors is inevitably limited, resulting in a significant reduction in amount of information available for matching [5, 6, 7]. Fig. 1 shows three representative challenging scenarios in partial fingerprint recognition, which can be even marked on fingers with large posture differences or poor quality (such as wear and wrinkles), leading to a substantial degradation in the performance of feature extraction and matching [8]. This not only reduces the user experience, but also poses great security risks [9].
@@
Figure 1: Partial fingerprint recognition faces three challenges: sparse minutiae distribution, similar local texture and significant modal difference. This figure shows corresponding representative examples: (a) Three weakly-textured fingerprints that lack appropriate landmark feature points. (b) Local ridge patterns of fingerprints from different fingers have high similarity. (c) Fingerprint features in different modalities (in this case from left to right are dry, normal and wet respectively) may be missed (green) or incorrectly extracted (red). Visualized feature points are minutiae extracted by VeriFinger [10].
@@
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [11, 12]. While certain studies use neural networks to extract region descriptors [13, 14, 15], consequently reducing the dependence on point features, these approaches require either a considerable effective area or exceedingly precise alignment. A series of fusion strategies are subsequently proposed at feature [16, 17, 18], image [19, 20] and decision levels [21, 22, 23], but most of them still depend on single-to-single matching or require several densely distributed high-quality images. In recent years, this field has garnered significant interest among researchers, leading to the development of several verification methods [24, 25, 8]. However, to the best of our knowledge, almost all existing fingerprint recoginition methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where joint decision is suggested about whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [26, 27, 28], in order to better balance robustness and speed in the overall process, rather than being made individually.
@@
Figure 2: Frameworks of different fingerprint recognition algorithms. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. The ‘Information Interaction’ extracts and compares features from paired data rather than executing independently and separately. It should be noted that the pose predictor in (a)(c)(d) and (b) estimate the relative and absolute pose respectively.
@@
In this paper, we proposed a multi-task CNN-Transformer hybrid network which jointly estimates identity verification and relative pose for partial fingerprints, called Partial Fingerprint Jigsaw Puzzle Network (FingerPuzzle). The intrinsic connection between feature correspondence and position correlation is exploited, aiming to promote each other to achieve higher performance. Naturally, the efficiency is also improved because these two tasks tasks that are independent in previous methods are compactly integrated. Inspired by works on natural images [29, 30, 31, 32, 33], we utilize a shared weight convolutional network to extract features of local patches and transformers to capture both local and global correlations. Besides, a fingerprint enhancement pre-training task is specially designed to enhance the compatibility with different image modalities and various texture patterns.
@@
Extensive experiments are conducted on abundant datasets, including several public datasets (NIST SD14 [34], FVC2002 DB1_A & DB3_A [35],FVC2004 DB1_A & DB2_A [36], FVC2006 DB1_A [37]) and an in-house dataset. Experimental results show that the proposed algorithm outperforms previous state-of-the-art methods in both partial fingerprint verification and pose estimation, and also demonstrates high efficiency.
@@
  • We propose a novel approach for partial fingerprint recognition by jointly estimating the identity verification and relative pose instead of the previous independent stages.
@@
  • Sufficient experiments are conducted on diverse datasets to comprehensively evaluate representative algorithms in partial fingerprints at different sizes, which demonstrates the superiority of our proposed method.
@@ II Related Work
According to different execution frameworks, existing single-to-single fingerprint recognition algorithms can be mainly divided into three categories: recognition based on key points, recognition based on region descriptors and verification based on pairwise correlation. We call the first two ‘recognition’ because highly generalized features are stored and compared here (in other words, each fingerprint only needs to extract features once) to enable efficient searches in large libraries. On the other hand, ‘verification’ uses richer information (features need to be re-extracted for each match) and is more suitable for confirmation scenarios with high precision requirements, such as unlocking a mobile phone. Fig. 2 shows the corresponding schematic diagrams of these processes.
II-A Recognition Based on Key Points
@@
Minutiae are the most popular features used for fingerprint matching [4]. Researchers define the features based on the attributes of minutiae themselves [11, 12] and auxiliary information (such as orientation [38], period map [39], ridge [40] etc.). Considering the lack of high-quality manually defined minutiae in latent fingerprints, Cao et al.[41] uniformly sampled images into small patches and used a network to extract fixed length vector as virtual minutiae. However, as explained in Fig. 1, these features are not discriminative enough in partial fingerprints n the other hand, some studies introduced universal key points in computer vision field into fingerprints [42, 7]. Features extracted in this way are more densely distributed, but at the same time more sensitive to details and therefore easily confused. There are also algorithms that exploit Level-3 features from high-resolution images [43, 44, 45], which are relatively expensive to deploy for civilian equipment. After getting satisfactory point sets, correspondence between them will be established based on the similarity of features and spatial structures, and then used to calculate relative poses and matching scores.
II-B Recognition Based on Region Representation
@@
Early studies use hand-crafted statistics to characterize fingerprint regions and convert them into fixed length vectors [46, 47, 48, 49]. This feature form significantly improves the speed of indexing at the expense of certain accuracy, and is usually used for rough classification. Engelsma et al.[13] used an end-to-end network to extract fixed-length global representations of minutiae and textures, which greatly improved the performance. Inspired by this, researchers have conducted more exploration and improvement of fixed-length representation methods based on deep learning [14, 50, 51, 15]. These algorithms usually require pose rectification in advance, such as adaptively adjusting the image geometry through a spatial transformation layer [13, 15], using a separate pose estimation network to rectify the absolute finger pose [50, 51], or exploiting the positioning relationships implicit in minutiae [14]. The error of pose estimation is tolerated to a certain extent because the extracted features are a generalization of regional characteristics. However, current methods struggle to address scenarios with incomplete or significantly different regions.
II-C Verification Based on Pairwise Correlation
@@
Compared with key points or fixed-length descriptors, gray-level information of pixels contain richer and more discriminatory information about local textures [52]. Traditional algorithms use image correlation for fine-grained verification, which exhibits a certain reliability [52, 53, 54]. Some works on fingerprint distortion further increase the tolerance of ridge misalignment [55, 56]. With the development of deep learning, researchers apply neural networks to directly extract and verify features between each other, significantly improving the performance and efficiency [57, 24, 25, 58]. He et al.[8] use multi-rotation and multi-size cropped image pairs as multi-inputs to help the network better understand scenarios with large spatial transformations or small overlapping areas. It should be noted that a significant difference between correlation based verification and region descriptor based recognition is that the relative pose of a certain fingerprint is rectified instead of the respective absolute pose. This is because features of such methods are compared more precisely with spatial correspondence thus have stricter requirements for alignment. Obviously, the verification performance will significantly declines with inaccurate pose estimation [8].
@@
Figure 3: An overview of FingerPuzzle. Bars are presented on the left to indicate each phase, where the color definition refers to Fig. 2. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. Dotted parts are only used to assist training and could be pruned in practical verification tasks. The outputs ‘p’ and ‘H’ of respective task heads correspond to the classification probability (whether the input fingerprints come from the same finger) and rigid transformation parameters (relative translation and rotation) respectively.
@@ III Method
In this paper, we propose a CNN-Transformer hybrid network that jointly estimates identity verification and relative pose of partial fingerprint pairs. Fig. 3 gives the complete flowchart of the proposed algorithm. The overall process can be divided into four phases, which are indicated by different color bars in schematic: feature extraction (red), information interaction (gold), identity verification (blue) and pose prediction (green). For an input fingerprint pair, our algorithm first extracts the respective features in parallel using convolutional blocks with shared weights (in stem and stages 1, 2). These two sets of features are then concatenated and fed to transformer blocks in stage 3 (along with some extra microprocessing) for sufficient information interaction both locally and globally. Finally, respective output layers utilize the aggregated information to predict classification probabilities for authentication (stage 4 - C and classifier) and regression values for relative poses (stage 4 - R and regressor). It should be noted that these two tasks are jointly estimated in our proposed network instead of being performed independently as previous methods because we hope to utilize the mutually beneficial coupling relationship between each other, inspired by works on jigsaw puzzles [26, 27, 28]. In particular, pose-related regression layers could be pruned to improve efficiency since only identity identification is necessary in most practices.
III-A Feature Extraction
@@
Some studies have proven that early stacked convolution blocks can significantly improve the stability and performance of subsequent vision transformers [59, 60]. Moreover, the hybrid strategy of using early CNN and subsequent transformer to aggregate local details and global information has achieved great success in multiple fields [29, 30, 31, 32, 33]. Inspired by these, we also use convolution blocks to parallelly extract independent features from each input partial fingerprint in the early stages of proposed network.
@@
Specifically, the feature extraction module sequentially contains: (1) a $3\times 3$ convolutional stem with stride $1$ (output $32$ channels); (2) stacked encoder blocks in stage 2 for preliminary extraction of robust features, which consisting of two Feature Extraction (FE) blocks and a convolutional layer with stride $2$ for downsampling while doubling the channels ( distinguished other convolutions by *); (3) stacked residual blocks in stage 3 to further enhance the expressive ability of neural network while mitigating gradient problems in deep layers. In order to improve robustness in different image qualities, we introduce FE block from a simple but efficient baseline for image restoration tasks [61], which is formally similar to the combination of Mobile Convolution (MBConv) block [62] and Multilayer Perceptron (MLP) block [63] with some minor adjustments. The details of corresponding architecture is presented on the left side of Fig. 3, where ‘DConv’ and ‘CA’ are depthwise convolution [62] and channel attention [64] respectively.
@@
Furthermore, we designed a pre-training task and corresponding data augmentation strategies to help the network better understand and refine the essential features of fingerprints. As shown in Fig. 4, the pre-training network consists only of FE block and convolution without additional complex components, arranged in a U-shape for fingerprint enhancement follows [23, 65]. For data preparation of inputs, we randomly crop small patches ($128\times 128$ in this paper) from high-quality rolled fingerprints as prototypes to ensure sufficient texture patterns at each position while accelerating the training process. The representative examples of corresponding data augmentation are given on the right of Fig. 4, including:
@@
These augmentation strategies are randomly selected and combined to generate the final input image. On the other hand, the corresponding binary fingerprint calculated by VeriFinger SDK 12.0 [10] is used as the target of this enhancement task. The parameters of stem and stage 1 are then loaded into corresponding modules of the main network.
Fig. 5 shows representative examples of several typical enhancement methods. It can be seen that our approach is smoother than methods based on grayscale mapping [67] or filtering [68]. In addition, our approach also shows certain robust performance when regional information is missing or wrong. This qualitative comparison strongly proves that our pre-training strategy can help the network understand more essential and complete features, and also avoids additional cumbersome enhancement steps. Further quantitative results are given in ablation experiments below.
Figure 4: Illustration of fingerprint enhancement task for pre-training. The network architecture is shown on the left, where solid and dotted boxes indicate the pre-trained parameters will (or not) be loaded into corresponding modules in Fig. 3. The right subfigure gives representative examples of inputs (original & augmented image) and targets (enhanced image) generation.
@@
III-B Information Interaction
Inspired by works on natural images [29, 30, 31, 32, 33], we utilize transformers to capture long-range dependencies within and across paired features, thereby ensuring sufficient information interaction. The paired features $f_{1}$ and $f_{2}$ are first biased with a sinusoidal positional encoding term [29], and then spatially flattened into 1D sequences and fed into stacked transformers with interleaved self- and cross- attention following [69, 33]. Let $x$ and $y$ denote the 2D position, $i$ and $d$ denote the current and total dimensions, the additional absolute position information is calculated as:

### v2

# Joint Identity Verification and Pose Alignment for Partial Fingerprints

@@ Abstract
Currently, portable electronic devices are becoming more and more popular. For lightweight considerations, their fingerprint recognition modules usually use limited-size sensors. However, partial fingerprints have few matchable features, especially when there are differences in finger pressing posture or image quality, which makes partial fingerprint verification challenging. Most existing methods regard fingerprint position rectification and identity verification as independent tasks, ignoring the coupling relationship between them — relative pose estimation typically relies on paired features as anchors, and authentication accuracy tends to improve with more precise pose alignment. Consequently, in this paper we propose a method for joint identity verification and pose alignment of partial fingerprint pairs, aiming to leverage their inherent correlation to improve each other. To achieve this, we propose a multi-task CNN (Convolutional Neural Network)-Transformer hybrid network, and design a pre-training task to enhance the feature extraction capability. Experiments on multiple public datasets (NIST SD14, FVC2002 DB1_A & DB3_A, FVC2004 DB1_A & DB2_A, FVC2006 DB1_A) and an in-house dataset demonstrate that our method achieves state-of-the-art performance in both partial fingerprint verification and relative pose estimation, while being more efficient than previous methods.
@@ I Introduction
Fingerprints are unique patterns composed of ridges and valleys on the finger surface. Due to its easy collection, high stability and strong recognizability, this biological characteristic exhibits significant application value. As early as the nineteenth century, researchers had conducted systematic studies on fingerprints [1, 2]. The reliability of fingerprint recognition gained official recognition in the twentieth century and found extensive application within the judicial sphere [2]. With the advancement of sensors and algorithms, fingerprint recognition technology has been swiftly applied in civilian and commercial fields, such as portable devices, financial services, and access control systems, providing people with convenience and security [3].
@@
Fueled by consumer demand and technological progress, the requirements towards integration, miniaturization, and portability has become prominent within the consumer electronics sector. This trend compels manufacturers to consistently reduce the size of sensors to accommodate progressively compact device architectures. Fingerprint sensors, currently one of the most popular biometric modules, have also introduced various miniaturized solutions for various applications [3]. However, the size of fingerprint sensors is inevitably limited, leading to a significant reduction in the information available for matching [4, 5, 6]. Fig. 1 shows three representative challenging scenarios in partial fingerprint recognition, which can be even more difficult on fingers with large posture differences or poor quality (such as wear and wrinkles), resulting in a sharp decline in performance of feature extraction and matching [7]. This not only reduces the user experience, but also poses great security risks [8].
@@
Figure 1: Partial fingerprint matching faces three challenges: sparse minutiae distribution, similar local texture and significant modal difference. This figure shows corresponding representative examples: (a) Three weakly-textured fingerprints that lack appropriate landmark feature points. (b) Local ridge patterns of fingerprints from different fingers have high similarity. (c) Fingerprint features in different modalities (in this case from left to right are dry, normal and wet respectively) may be missed (green) or incorrectly extracted (red). Visualized feature points are minutiae extracted by VeriFinger [9].
@@
A natural idea for partial fingerprint matching is to transfer existing solutions for rolled or plain fingerprint recognition. Nevertheless, conventional algorithms usually rely on a sufficient number of key points to establish and compare spatial structural relationships [10, 11]. While certain studies use neural networks to extract fixed-length region descriptors [12, 13, 14, 15], consequently reducing the dependence on point features, these approaches require either large effective area or exceedingly precise alignment. A series of fusion strategies are subsequently proposed at feature [12, 16], image [17, 18] and decision levels [19, 20], but most of them still depend on one-to-one matching or require several densely distributed high-quality images. In summary, although these highly generalized feature forms ensure efficient matching, some local details are missed, which is not enough to support high-precision recognition for partial fingerprints due to their extremely small effective area. Considering this, some researchers simultaneously input fingerprint images and extract interrelated features to fully utilize complete information [7, 21], which effectively improves the matching performance on partial fingerprints. However, to the best of our knowledge, almost all existing fingerprint matching methods (compared in Fig. 2) regard pose rectification and identity verification as independent tasks, ignoring the coupling relationship between them. Specifically, the estimation of relative pose typically relies on paired regional features as anchors, and the accuracy of authentication generally increases with enhanced precision in pose alignment. This is similar to a pictorial jigsaw puzzle where joint decision is suggested about whether (are they neighboring image fragments? ) and how (what’s the reassemble solution? ) any two pieces fit together [22, 23], in order to better balance robustness and speed in the overall process, rather than being made individually.
@@
Figure 2: Frameworks of different fingerprint matching algorithms. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed independently. Feature Extraction in red and gold extract independent and interrelated features from paired data respectively. Pose Predictor in (a)(c)(d) and (b) estimate relative and absolute pose respectively.
@@
In this paper, we propose a multi-task CNN-Transformer hybrid network which Jointly performs Identity verification and Pose alignment for partial fingerprints, called JIPNet. Paired images are input instead of their simplified features to provide more complete information, enabling more flexible and precise analysis. The intrinsic connection between feature correspondence and position correlation is exploited, aiming to promote each other to achieve higher performance. Naturally, the efficiency is also improved because these two tasks are compactly integrated. Inspired by works on natural images [24, 25, 26, 27], we utilize a shared weight convolution layers to extract regional features and transformers to capture both local and global correlations. Besides, a fingerprint enhancement pre-training task is specially designed to enhance the compatibility with different image modalities and various texture patterns.
@@
Extensive experiments are conducted on multiple datasets, including several public datasets (NIST SD14 [28], FVC2002 DB1_A & DB3_A [29], FVC2004 DB1_A & DB2_A [30], FVC2006 DB1_A [31]) and an in-house dataset. Experimental results show that the proposed algorithm outperforms previous state-of-the-art methods in both partial fingerprint verification and pose estimation, and also demonstrates high efficiency.
@@
  • We propose a novel approach for partial fingerprint recognition by jointly estimating the authentication probability and relative pose instead of the previous independent stages.
@@
  • Extensive experiments are conducted on diverse datasets to evaluate representative state-of-the-art algorithms, which demonstrates the superiority of our proposed method.
@@ II Related Work
According to different execution frameworks and feature representation forms, existing one-to-one fingerprint matching algorithms can be mainly divided into three categories: key points based, fixed-length region descriptor based, and pairwise interrelated feature based. In the first two types of algorithms highly generalized features are stored and compared here (in other words, each fingerprint undergoes feature extraction only once) to enable efficient searches in large-scaled galleries. On the other hand, pairwise interrelated feature based methods could use richer information (features need to be re-extracted for each match) and is more suitable for confirmation scenarios with high precision requirements and few gallery / reference fingerprints, such as unlocking a mobile phone. Given the multitude of fingerprint matching algorithms, this rough classification may not comprehensively encompass all works. Nonetheless, we have endeavored to summarize them to the best of our knowledge. Fig. 2 shows the corresponding schematic diagrams of these processes.
II-A Key Points Based
@@
Minutiae are the most popular features used for fingerprint matching [3]. Researchers define minutiae descriptors using the attributes of minutiae themselves [10] and auxiliary information (such as orientation [32], period map [33], ridge [11] etc.). Considering the lack of reliable minutiae in latent fingerprints, Cao et al.[19] uniformly sampled images into small patches and used a network to extract fixed length vector to describe virtual minutiae. However, as explained in Fig. 1, these features are not discriminative enough in partial fingerprints. On the other hand, some studies introduced universal key points in computer vision field into fingerprints [34, 6]. Features extracted in this way are more densely distributed, but at the same time more sensitive to unreliable image details and therefore easily confused. There are also algorithms that exploit Level-3 features from high-resolution images [35, 36], which are relatively expensive to deploy in most commercial applications. After getting key point sets, correspondence between them will be established based on the similarity of features and spatial structures, and then relative poses and matching scores are calculated.
II-B Fixed-length Region Representation Based
@@
Early studies use hand-crafted statistics to characterize fingerprint regions and convert them into fixed length vectors [37, 38, 39]. This feature form significantly improves the speed of large database indexing at the expense of reduced accuracy, and is usually used for rough searching. Engelsma et al.[12] propose an end-to-end network to extract fixed-length global representations of minutiae and textures, which greatly improves the performance of fixed-length representation methods. Inspired by this, researchers have conducted more exploration and improvement of fixed-length representation methods based on deep learning [14, 40, 15]. In addition, Gu et al.[13] conducted dense sampling on complete fingerprints and compared the fixed length descriptors of each patch, ultimately making a comprehensive decision. These algorithms usually require pose rectification in advance, such as adaptively adjusting images through a spatial transformation layer [12, 15], rectifying absolute pose of fingerprints through a separate network [40], or exploiting the positioning relationships implicit in minutiae [14]. The error of pose estimation is tolerated to a certain extent because the extracted features are a generalization of regional characteristics. However, current methods may struggle to address partial fingerprints, as the pose estimation on them is often unstable.
II-C Pairwise Interrelated Feature Based
@@
Compared with key points or fixed-length descriptors, gray-level features of pixels contain richer and more discriminatory information about local textures [41]. Traditional algorithms use image correlation for fingerprint verification, which can be affected by skin distortion [41, 42]. Some works on fingerprint distortion further increase the tolerance of ridge misalignment [43]. With the development of deep learning, researchers apply neural networks to directly extract and verify interrelated features between each other, significantly improving the verification performance and efficiency [44, 45, 46]. He et al.[7] use multi-rotation and multi-size cropped image pairs as combined inputs to help the network better understand scenarios with large spatial transformations or small overlapping areas. It should be noted that a significant difference between this type of method and methods based on fixed-length region descriptor is that the relative pose of a certain fingerprint is rectified instead of the respective absolute pose. Features of such methods are then compared more precisely with spatial correspondence thus have stricter requirements for alignment. Obviously, the verification performance will significantly decline with inaccurate pose estimation [7].
@@
Figure 3: An overview of JIPNet. Bars are presented on the left to indicate each phase, where the color definition refers to Fig. 2. The process passed by parallel arrows indicates that the corresponding modules share weights and their functions can be executed in parallel. * represents the number of channels are doubled after the corresponding convolution. Dotted parts are only used to assist training and could be pruned in practical verification tasks. The outputs ‘p’ and ‘H’ of respective task heads correspond to the classification probability (whether the input fingerprints come from the same finger) and rigid transformation parameters (relative translation and rotation) respectively.
@@ III Method
In this paper, we propose JIPNet, a CNN-Transformer hybrid network, to jointly estimates the authentication probability and relative pose of partial fingerprint pairs. Fig. 3 gives the complete flowchart of our proposed algorithm. The overall process can be divided into four phases, which are indicated by different color bars in schematic: independent feature extraction (red), interrelated feature extraction (gold), identity verification (blue) and pose prediction (green). For an input fingerprint pair, our algorithm first extracts the respective features in parallel using convolutional blocks with shared weights (in Stem and Stages 1, 2). These two sets of features are then concatenated and fed to transformer blocks in Stage 3 (along with some extra microprocessing) for sufficient information interaction both locally and globally. Finally, respective output layers utilize the aggregated information to predict classification probabilities for authentication (Stage 4 - C and Classifier) and regression values for relative poses (Stage 4 - R and Regressor). It should be noted that these two tasks are estimated in an integrated manner instead of being performed independently as previous methods because we hope to utilize the mutually beneficial coupling relationship between each other, inspired by works on jigsaw puzzles [22, 23]. In particular, pose-related regression layers could be pruned to improve efficiency since only identity identification is necessary in most applications.
III-A Independent Feature Extraction
@@
Previous studies have proven that early stacked convolution blocks can significantly improve the stability and performance of subsequent vision transformers [47, 27]. Similarly, we also use convolution blocks to parallelly extract independent features from each input partial fingerprint in the early stages of proposed network.
@@
Specifically, the feature extraction module sequentially contains: (1) a $3\times 3$ convolutional stem with stride $1$ (output $32$ channels); (2) stacked encoder blocks in Stage 2 for preliminary extraction of robust features, which consisting of two Feature Extraction (FE) blocks and a convolutional layer with stride $2$ for downsampling; (3) stacked residual blocks in Stage 3 to further enhance the expressive ability of neural network while mitigating gradient problems in deep layers. Among them, FE block is introduced from a simple but efficient baseline for image restoration tasks [48], which is formally similar to the combination of Mobile Convolution (MBConv) block [49] and Multilayer Perceptron (MLP) block [50], to improve the robustness of feature encoding in different image qualities. The details of corresponding architecture is presented on the left side of Fig. 3, where ‘DConv’ and ‘CA’ are depthwise convolution [49] and channel attention [51] respectively.
@@
Furthermore, we design a pre-training task and corresponding data augmentation strategies to help the network better extract and reconstruct the essential features of fingerprints. As shown in Fig. 4, the pre-training network consists only of FE block and convolution without extra complex components, arranged in a U-shape for fingerprint enhancement follows [20, 52]. For data preparation of inputs, we randomly crop small patches ($128\times 128$ in this paper) from high-quality rolled fingerprints as prototypes to ensure sufficient texture patterns at each position while accelerating the training process. Corresponding cases after augmentation are given on the right of Fig. 4, including:
@@
These augmentation strategies are randomly selected and combined to generate the final input image. On the other hand, the corresponding binary fingerprint calculated by VeriFinger SDK 12.0 [9] is used as the enhancement target. Parameters of Stem and Stage 1 are then loaded into corresponding modules of JIPNet.
Fig. 5 shows representative examples of several typical enhancement methods. It can be seen that our approach is smoother than methods based on grayscale mapping [54] or filtering [55]. In addition, our approach also shows certain robust performance when regional information is missing or wrong. This qualitative comparison strongly proves that our pre-training strategy could help the network understand more essential and complete features, and also avoids additional cumbersome enhancement steps. Further quantitative results are given in ablation experiments below.
Figure 4: Illustration of fingerprint enhancement task for pre-training. The network architecture is shown on the left, where solid and dotted boxes indicate the pre-trained parameters will (or not) be loaded into corresponding modules in JIPNet. The right subfigure gives representative examples of inputs (original & augmented image) and targets (enhanced image) generation.
@@
III-B Interrelated Feature Extraction
The hybrid strategy of using early CNN and subsequent transformer to aggregate local details and global information has achieved great success in multiple fields [24, 25, 26]. Inspired by these, we utilize transformers to capture long-range dependencies within and across paired features, thereby ensuring sufficient information interaction. The paired features $f_{1}$ and $f_{2}$ are first biased with a sinusoidal positional encoding term [24], and then spatially flattened into 1D sequences and fed into stacked transformers with interleaved self- and cross- attention. Let $x$ and $y$ denote the 2D position, $i$ and $d$ denote the current and total dimensions, the additional absolute position information is calculated as: